Been trying all night to figure this out with my child....can't seem to find a pattern
101,316,192,225,283...
just need to find the next number.
if you can help please show

Found 2 solutions by LinnW, Theo:
Answer by LinnW(1048)

Notice the progression for the 1st, 3rd and 5th elements
101 192 and 283
192 is 91 more than 101
283 is 91 more than 192
Now look at the 2nd, 4th elements
316 , 225
The difference for these two is 91,
so one could argue that the 6th element should
be 225 - 91 = 134
and the whole sequence would be
101,316,192,225,283,134

Answer by Theo(13342)

if i skip numbers, then i get + 91 for the odd numbers terms and -91 for the even numbered terms.

192 - 101 = 91
283 - 192 = 91

that's the odd numbered terms.

225 - 316 = -91 for the even numbered terms.

if i assume that's a valid rule, then the next term in the sequence is an even numbered term and would be equal to 134 because 225 - 91 = 134.

i don't see any other patterns, but i'm not the expert in dealing with these type problems, so that doesn't mean there isn't any.

this is my guess, anyway.